OpenStudy (anonymous):

the position of a particle as it moves along the axis is given for t>0 by x=(t^3-3t^2+6t)m,where t is in s.where is the particle when it achieves its minimum speed (after t=0)

OpenStudy (anonymous):

@zepdrix @Miracrown

OpenStudy (anonymous):

i think u have to do double differentiation here ...

OpenStudy (anonymous):

@HS_CA @CallMeKiki @countrygirl1431 @carolinar7 @Camila1315 @Ivanskodje @isry98 @just_one_last_goodbye @kobi @Loser66 @lordhelix8th @LoneWolfKay

OpenStudy (anonymous):

@SolomonZelman @ParthKohli

OpenStudy (anonymous):

@Nnesha @Abhisar

OpenStudy (abhisar):

Do you know about minima and maxima?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

if we compute the first derivative of your function \(x(t)\), we get: \[v(t)=\frac{{dx}}{{dt}} = 3{t^2} - 6t + 6\] which is the equation of a parabola. Now the minimum speed occurs at the t-coordinate of the vertex of such parabola. Such value of \(t\) is: \[t_0 = - \frac{{ - 6}}{{2 \cdot 3}} = ...?\] now, please replace \(t=1\) into formula of \(x(t)\), or in other words, compute the quantity \(x(t_0)\)